Circular No. 7673

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2001 O2 K. J. Lawrence, S. Pravdo, and E. F. Helin, Jet Propulsion Labora-

tory, report the discovery by NEAT of a diffuse comet with some central condensation; on July 29, it showed nebulosity 12′′ toward the east.

2001 UT  α_{2000}  δ_{2000}  m_{1}
July 25.42122  22^{h}44^{m}33.^{s}55  −24^{o}01′37′′.9  19.3
25.43195  22 44 33.15  −24 01 36.8  19.4
25.44270  22 44 32.74  −24 01 37.3  19.4
29.37154  22 41 48.73  −24 02 58.3  18.6
29.38217  22 41 48.33  −24 02 58.6  18.1
29.39278  22 41 47.88  −24 02 58.8  19.2

SUPERNOVA 2001di IN UGC 3259 G. M. Hurst, Basingstoke, England, reports the discovery of an appar-

ent supernova (mag 17.7) by M. Armstrong (Rolvenden, England) on an unfiltered CCD image (limiting mag ∼ 19) secured for the U.K. Nova/Supernova Patrol on July 28.090 UT with a 0.36-m Schmidt-Cassegrain telescope. A follow-up series of images was obtained on July 30.07 with thin cloud present (with the new object still at mag ∼ 17.7). SN 2001di is located at α = 5^{h}14^{m}17^{s}.57, δ = +72^{o}20′41′′.0 (equinox 2000.0), which is 6′′ east and 39′′ north of the nucleus of UGC 3259. A CCD image (limiting mag 19.5) taken by Armstrong on 2000 Nov. 3 shows nothing at this position, and nothing is present on the following Palomar Sky Survey images: 1953 Oct. 14, 1991 Dec. 7 (limiting blue mag 22.5), 1996 Oct. 2 (infrared).

V1548 AQUILAE D. Lynch, R. J. Rudy, S. Mazuk, and C. Venturini, Aerospace Cor-

poration; and R. C. Puetter, Center for Astrophysics and Space Science, University of California, San Diego, report 0.8–2.5-μm spectrophotometry of V1548 Aql with the Lick Observatory 3-m telescope (+ Aerospace Near Infrared Imaging Spectrograph) on July 8.42 UT, ∼ 55 days after discovery: “The spectrum indicated fairly low excitation, showing the fluorescently excited O I lines and the Fe II ‘1-μm’ lines (Rudy et al. 2000 Ap.J. 539, 166), as well as permitted lines of H I and He I. Both the He I 1.0830-μm triplet and He I 2.0581-μm singlet displayed P-Cyg profiles. All the line widths were ∼ 1000 km/s. The continuum decreased monotonically with increasing wavelength and showed no evidence of thermal emission from dust.”

2001 July 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green